OpenStudy (anonymous):

how do you point 2y-3x=1 on a slope graph

OpenStudy (anonymous):

let x=0 find value of y------>got the point (0,y) let y=0 find the value of x----->'' "(x,0)

OpenStudy (anonymous):

first find y in function of x : \[y = \frac{ 1+3x }{ 2 }\] so you got a line in the form \[y = ax +b\] \[a\] is the slope (3/2) and \[ b\] is the linear coefficient (1/2) with x = 0 you got y = 1/2 thats where the line crossed the y axis and with y = 0 you have \[\frac{ 1 }{ 2 } + \frac{ 3x }{ 2 } = 0\]x = -1/3( where the line crosses the x axis and you can plot it now.

OpenStudy (anonymous):

thanks so much... i have been working on this promble for a hour